# Configurations, Tessellations and Tone Networks

### v4

@@ V Pentatonic Tone Network
Figure 12: The pentatonic tonnetz is represented here in two isomorphic ways as the Levi graph of the Desargues configuration. Each two-note chord belongs to three distinct three-note chords and each three-note chord contains three distinct two-note chords. The girth of the graph is six, pointing to the existence of hexacycles, for example $\langle CDA,\,CD,\,CDG,\,CG,\,CGA,\,CA,\,CDA\rangle$. The graph is manifestly self-dual. One also sees that there is a Hamiltonian cycle embracing all twenty major and minor chords. On the right, twenty hexacycles are clearly visible.
Looking at the Desargues configuration in Figures 10 and 11 one might be tempted to assign some special significance to the point of perspective $V$ or in the musical picture the chord $GA$. But this is illusory – any of the ten points can be taken as a point of perspective, from which the two associated triangles can be identified and the line of perspective.  [^17] Nonetheless, once the symmetry is broken by the choice of a point of perspective the other points line up to it in a specific way with certain relations among them. Thus, the choice of a preferred chord, such as $GA$, as the point of perspective is analogous to a choice of key. No one choice is preferred a priori, but once a choice has been made then the music associated with that chord involves a specific set of relations between that chord and the other chords. This is similar to the way in which in common practice tonality no specific tonal key is preferred a priori to another – they are all on an equal footing – but once a key has been chosen then much else follows. Thus, the two-letter ‘major’ chords determine ten major keys; and since the configuration is self-dual, the ten three-letter ‘minor’ chords determine ten minor keys. A more symmetrical way of drawing the Desargues configuration can be found in figure 2.3a of Coxeter (1974) in which two different points of perspective, and hence two different lines of perspective as well, are treated in a manifestly symmetric way.
In summary, we have shown the following:
@@
Can such a setup be generalized? It should be apparent that when we have an odd number of basic tones – for example, any system of music based on an odd number of tones per octave, with or without equal temperament – the geometric combinatorics of the pentatonic system can be generalized as follows:
@@
VI Subset relations vs pitch relations
A possible criticism of our approach to the construction of a tonnetz for the pentatonic system (and for the more general systems of tonnetze considered in Proposition 4) might be that it is based on subset inclusion relations rather than pitch relations along the lines of the $LPR$ operations of neo-Riemannian theory that form the basis of the Eulerian tonnetz. Prima facie, these do seem to be very different constructions. Nonetheless, it has to be admitted that the Levi graphs of the Eulerian configuration and the pentatonic configuration bear a striking ‘family resemblance’, suggesting some commonality in the superficially divergent methodologies we have employed. So what is going on?
@@ VI Subset relations vs pitch relations
The pentatonic system, which we have put forth on the basis of set inclusion relations, is evidently not quite rich enough to allow for much discussion about pitch relations and in any case the ‘chords’ have different cardinalities. But what can be said about subset relations in the case of the Eulerian tonnetz? Here, it turns out that some interesting observations can be made. Suppose we consider the twelve notes of the chromatic scale and the twelve major chords. Clearly, each chord contains three notes; but a little reflection shows also that each note is contained in exactly three major chords. For example, the note $C$ is contained in $CM$, FMFM and $A_{\flat}M$. So, in fact, we have a three-to-three map between the note set and the major chord set. It is also true that any note is contained in exactly three minor chords, so one obtains another three-to-three map. What do these maps look like? In each case we can construct the relevant Levi graphs, and the results are shown in Figure 13. Curiously, each of these graphs has the incidence structure of the Eulerian tonnetz. So evidently the combinatorial structure of the tonnetz is already implicit in the relation between the major chords and the tones within these chords. More precisely, we have the following:
@@
The Eulerian tonnetz can be constructed purely on the basis of set inclusion relations, without reference to voice leading, minimal work or efficiency.
@@
Begin with the twelve notes of chromatic scale and the twelve major chords. Associated with each note are the three major chords to which it belongs. For example, $C$ belongs to $CM$, $A_{\flat}M$ and $FM$. One can then construct a hexagonal tessellation of the plane in which each note is surrounded in a symmetric way by its three associated major chords. Each hexagon in the tessellation then consists of a cycle of notes and major chords. For example, the note $C$ is contained in the cycle $\langle C,A_{\flat}M,E_{\flat},E_{\flat}M,G,CM,C\rangle$. The three notes in any given cycle can be combined to form a minor chord, so we mark that the center of that hexacycle with that minor chord. For example, the cycle just considered is marked with $Cm$ at the center. Each minor chord is then connected to the three notes that it contains and to the three major chords in the hexagon at the center of which the minor chord sits. The resulting triangular tessellation of the plane is the grand Eulerian tonnetz shown in Figure 14. If then we rub out the notes and the links between notes and chords, we are left with the Eulerian tonnetz as it appears in Figure 1. ∎
@@
Figure 14: The complete triadic tonnetz is a face-centered hexagonal tessellation admitting an infinite tripartite graph based on an underlying triangular tessellation. This infinite collection of face-centered hexagons, each containing a pitch at the center, has the property that each major chord is surrounded its three constituent pitches and each minor chord is surrounded by its three constituent three pitches. If the pitches and the edges joining them are rubbed out, then one is left with the major-minor tessellation of Figure 1; and then if the enharmonic-equivalent pitches are identified, one is led to Figure 2. If one or the other of the remaining classes of edges is rubbed out instead, by the same procedure one obtains the Levi graphs of Figure 14.
@@
This may come as a surprise, but the matter can be understood readily enough. It is well-known that the tonnetz has two different ‘chicken-wire’ representations, one in which the notes are at the vertices and one in which the major and minor chords are at the vertices. The note-based version of tonnetz is essentially that of Euler and has been used by Cohn (2012) and others as a starting point for musical analyses. Chord-based versions of the tonnetz can be found in Waller (1978), Douthett & Steinbach (1998) and many other references. We say that a major chord and a minor chord are incident if they share two notes in common. This relation also functions as a basis for voice leading. The point is that one need not start from that relation as the primary one.
@@ VII Remarks on the Twelve-Tone System
The role of unordered hexachords in twelve-tone music has long been recognized. For example, Lewin (1968) approaches the structure of Schönberg’s Violin Fantasy with a study based on the unordered hexachord $F,G,A,B_{\flat},B,C_{\sharp}]$ and its complement $C,D,D_{\sharp},E,F_{\sharp},G_{\sharp}]$. Unordered hexachords and their complements also played a prominent role in the work of Josef Matthias Hauer, whose techniques anticipated those of Schönberg. It is not unreasonable then to take the view that at the very outset of any approach to the theory of twelve-tone music the combinatoric geometry of such unordered sets should be investigated in the spirit we have pursued the Eulerian and pentatonic tonnetze.
@@
At first glance, such an assignment might seem structureless; but the permutation group for six objects is very special and the highly nontrivial structure of this group is directly related to the particular type of tone network that we wish to propose for twelve-tone music. The method that we consider is based on Sylvester’s theory of duads and synthemes (Sylvester 1844, 1861). A duad is an unordered pair of distinct tones taken from $H_{0}$. A little thought shows there are fifteen duads. Thus, a duad $D$ is an unordered set of the form
@@
Thus, we have constructed a combinatoric configuration. More precisely, we have two sets, viz., the set of duads and the set of synthemes.
@@
Table 1: Each of the six letters corresponds to a collection of five non-overlapping number-synthemes called a number-total. In each number-total one finds all fifteen of the number-duads.
@@ 66 $af,bc,de$ $ad,be,cf$ $ae,bf,cd$ $ac,bd,ef$ $ab,ce,df$
Table 3: Each of the six numbers corresponds to a collection of five non-overlapping letter-synthemes called a letter-total. In each letter-total one finds all fifteen of the letter-duads.
@@
Each has fifteen elements and they are in three-to-three correspondence. We say that the duad $D$ is incident with the syntheme $S$ if it is among the three duads contained in $S$. Inspection of Tables II and IV shows that two synthemes have at most one duad in common. This gives us an incidence structure and we can construct the corresponding Levi graph.
In fact, there is one more layer of structure, which comes as a bit of a surprise, and is of importance, and this is the following. Suppose we consider a collection of five synthemes with the property that each duad appears exactly once. This makes sense since there are fifteen duads and a collection of five synthemes must include fifteen duads. Such a collection of five ‘non-overlapping’ synthemes is called a ‘total’ (again, Sylvester’s term). It turns out that there are exactly six such totals that can be constructed.
@@
Let us label the six totals with lower case Roman letters. Thus, the totals are elements of the set $\{a,b,c,d,e,f\}$. Since we have a second set of six elements, we can form the corresponding ‘duads’ and ‘synthemes’ for this set as well. We shall call these letter-duads and letter-synthemes, respectively. Thus, a letter-syntheme is an unordered triple of letter-duads. Then by a letter-total we mean a collection of five non-overlapping letter-synthemes. Two distinct totals have one syntheme in common. Hence, the letter duads of the totals are the original synthemes; and it follows that the letter-synthemes are the original duads. The relations between duads, synthemes, totals, letter-duads, letter-synthemes, and letter-totals can be seen in Tables I – IV (Richmond 1900, Baker 1925, Coxeter 1958).
@@
The advantage of using the number-duads and the letter-duads is that this allows us to label the resulting Levi graph neatly, where the number-duads, which correspond to letter-synthemes, are represented by white vertices and the letter-duads, which correspond to synthemes, are represented by black vertices. In summary, one sees that there is indeed a rich hierarchical structure implicit in the specification of a hexachord.
@@
The Levi graph of the Cremona-Richmond $\{15_{3}\}$ is shown in Figure 15 (Tutte 1947, Coxeter 1950, Brier & Bryant 2021). We can take this graph as representing a tonnetz for the twelve-tone system when an unordered hexachord has been selected as a primitive. Its girth is eight: there are no hexacycles.  [^18] A construction of the $\{15_{3}\}$ is as follows. We begin with six points $p_{1},.\,.\,.,p_{6}$ in general position in $\mathbb{RP}^{4}$. When joined pairwise, these points determine fifteen lines. Let us write $L_{12}=p_{1}\vee p_{2}$ for the line joining $p_{1}$, $p_{2}$, and $H_{3456}=\vee(p_{3}p_{4}p_{5}p_{6})$ for the complementary hyperplane joining $p_{3}$, $p_{4}$, $p_{5}$, $p_{6}$. This hyperplane is an $\mathbb{RP}^{3}$. Each of the fifteen lines meets its complementary hyperplane at a point. Thus, for example, we can write $P_{12}=L_{12}\wedge H_{3456}$ for the meet of $L_{12}$ and the hyperplane $H_{3456}$. There are evidently fifteen such ‘meeting points’. On the other hand, for example, we also have $P_{34}=L_{34}\wedge H_{1256}$ and $P_{56}=L_{56}\wedge H_{1234}$. Let us write $S_{12,34,56}$ for the line given by the meet of the three hyperplanes. Since $H_{1234}$ and $H_{1256}$ both contain $L_{12}$ it follows that the plane of intersection $D_{12}$ of these two hyperplanes contains $L_{12}$. Then since $D_{12}$ meets $H_{3456}$ at $S_{12,34,56}$ it must be that $L_{12}$ meets $S_{12,34,56}$ at $P_{12}$. By symmetry, all three of the ‘duad’ meeting points $P_{12}$, $P_{34}$, $P_{56}$ lie on the ‘syntheme’ line $S_{12,34,56}$. The fifteen meeting points are the duads of the configuration; and the fifteen lines of three-at-a-time hyperplane intersection, each of which contains three of the duads, are the synthemes of the configuration. That gives the four dimensional version of the construction, which can be projected to a plane to give a $\{15_{3}\}$ in two dimensions, the Cremona-Richmond configuration.
Thus, summing up the results of this section, we have the following:
@@
A tonnetz can be associated with the selection of an unordered hexachord in $12$-tone music, given by a self-dual $\{15_{3}\}$ in $\mathbb{R}^{2}$ known as the Cremona-Richmond configuration. In this tonnetz, the $15$ ‘major’ duad chords are represented by points and the $15$ ‘minor’ syntheme chords are represented by lines. The $45$ incidence relations between the $15$ points and the $15$ lines determine the edges of the corresponding Levi graph, which is known as Tutte’s $8$-cage.
Figure 15: The Levi graph of the Cremona-Richmond configuration $15_{3}$, also known as Tutte’s 88-cage, maps to a tonnetz for the system of twelve-tone music based on an unordered hexachord. Each white vertex represents a duad (number pair) and each black vertex represents a syntheme (letter pair). Each duad belongs to three synthemes and each syntheme contains three duads.

### v5

@@ V Pentatonic Tone Network
Figure 12: The pentatonic tonnetz is represented here in two isomorphic ways as the Levi graph of the Desargues configuration. Each two-note chord belongs to three distinct three-note chords and each three-note chord contains three distinct two-note chords. The girth of the graph is six, pointing to the existence of hexacycles, for example $\langle CDA,\,CD,\,CDG,\,CG,\,CGA,\,CA,\,CDA\rangle$. The graph is manifestly self-dual. One also sees that there is a perimeter Hamiltonian cycle embracing all twenty major and minor chords. In fact, there are twenty-four such icosacycles altogether. On the right, the twenty hexacycles of the pentatonic tonnetz are clearly visible.
Figure 13: A simple composition, ‘On the Perimeter’ based on a Hamiltonian cycle of the pentatonic tonnetz, using a Japanese pentatonic scale with the five notes $C$, $D$, $E_{\flat}$, $G$, $A_{\flat}$. The Hamiltonian 0pp-icosacycle is represented in this piece by the five overlapping 1pp-hexcycles of the pentatonic tonnetz, each played twice. See the Appendix, Table VII, for a tabulation of the cycles of the Desargues tonnetz.
@@
For any system of music with an odd number of tones per octave, with or without equal temperament, the combinatorics of the pentatonic system can be extended as follows:
@@
VI Subset relations vs voice-leading relations
A possible criticism of our approach to the construction of a tonnetz for the pentatonic system (and for the more general systems of tonnetze considered in Proposition 4) might be that it is based on subset inclusion relations rather than voice-leading relations along the lines of the familiar $L$, $P$ and $R$ operations of neo-Riemannian theory that form the basis of the Eulerian tonnetz.
Prima facie, these do seem to be very different constructions. Nonetheless, it has to be admitted that the Levi graphs of the Eulerian configuration and the pentatonic configuration bear a striking ‘family resemblance’, suggesting some commonality in the superficially divergent methodologies we have employed.
@@ VI Subset relations vs voice-leading relations
This suggests that we should look at subset relations in the Eulerian tonnetz. In this respect, some interesting observations can be made. Consider the twelve notes of the chromatic scale and the twelve major chords. Clearly, each chord contains three notes; but it is also the case that each note is contained in exactly three major chords. For example, the note $C$ is contained in $CM$, FMFM and $A_{\flat}M$. So, we have a three-to-three map between the note set and the major chord set. It is true as well that any note is contained in exactly three minor chords – so one obtains another three-to-three map. What do these maps look like? We can construct the relevant Levi graphs, and the results are shown in Figure 14. Curiously, each of these graphs has the incidence structure of the Eulerian tonnetz. So, evidently, the combinatorial structure of the tonnetz is already implicit in the relation between the major chords and the tones within these chords. More precisely, we have the following:
@@
The Eulerian tonnetz can be constructed purely on the basis of set inclusion relations, without reference to voice-leading relations, minimal work or parsimony.
@@
Begin with the twelve notes of chromatic scale and the twelve major chords. Associated with each note are the three major chords to which it belongs. For example, the note $C$ belongs to the chords $CM$, $A_{\flat}M$ and $FM$. One can then construct a hexagonal tessellation of the plane in which each note is surrounded in a symmetric way by its three associated major chords. Each hexagon in the tessellation then consists of a cycle of notes and major chords. For example, the note $C$ is contained in the cycle $\langle C,A_{\flat}M,E_{\flat},E_{\flat}M,G,CM,C\rangle$. Now, the three notes in any given cycle can be combined to form a minor chord, so we mark that the center of that hexacycle with that minor chord. For instance, the cycle just considered is marked with $Cm$ at the center. Each minor chord is then connected to the three notes that it contains and to the three major chords in the hexagon at the center of which the given minor chord sits. The resulting triangular tessellation of the plane is the complete triadic tonnetz shown in Figure 15. If we rub out the notes and the links between notes and chords, we are left with the Eulerian tonnetz of Figure 1. ∎
@@
Figure 15: The complete triadic tonnetz is a face-centered hexagonal tessellation admitting an infinite tripartite graph based on an underlying triangular tessellation. This infinite collection of face-centered hexagons, each containing a pitch at the center, has the property that each major chord is surrounded its three constituent pitches and each minor chord is surrounded by its three constituent pitches. If the pitches and the edges joining them are rubbed out, then one is left with the major-minor tessellation of Figure 1; and then if the enharmonic-equivalent pitches are identified, one is led to Figure 2. If one or the other of the remaining classes of edges is rubbed out instead, then by the same procedure one obtains the Levi graphs of Figure 14.
The idea that the Eulerian tonnetz can be constructed on the basis of set-inclusion relations may come as a surprise, but the matter can be understood readily enough. It is well-known that the tonnetz has two different ‘chicken-wire’ representations, one in which the notes are at the vertices and one in which the major and minor chords are at the vertices.
@@
The note-based version of tonnetz is essentially that of Euler and has been used by Cohn (2012) and many others as a starting point for musical analyses. Chord-based versions of the tonnetz can be found in Waller (1978), Douthett & Steinbach (1998), Tymoczko (2011b) and other references. We say that a major chord and a minor chord are incident if they share two notes in common. This relation also functions as a basis for voice leading. The point is that one need not start from that relation as the primary one. Alternatively, if one takes the note/major-chord tonnetz of Figure 14 as given, then we observe that the minor chords are in one-to-one correspondence with the 2pp-hexacycles of this graph. The three minor chords to which a note belongs are the three 2pp-hexacycles that contain that note.
One sees, for example, that the note $C$ is contained in the 2pp-hexacycles corresponding to $Cm$, AmAm and $Fm$. This point of view is interesting since it breaks the duality between the major and minor chords that seems to dominate the usual philosophy of the tonnetz and to some extent restores the primacy of the major chords. According to this view, the minor chords can be identified as secondary structures – namely, the 2pp-hexacycles. In that case, one can also ask about the role of the four 3pp-hexacycles of the note/major-chord tonnetz. Inspection of the Levi graph shows that these correspond to the four augmented triads.
@@ VII Remarks on the Twelve-Tone System
The role of unordered hexachords in twelve-tone music has long been recognized. For example, Lewin (1968) approaches the structure of Schönberg’s Violin Fantasy with a study based on the unordered hexachord $F,G,A,B_{\flat},B,C_{\sharp}]$ and its complement $C,D,D_{\sharp},E,F_{\sharp},G_{\sharp}]$.
Unordered hexachords and their complements also played a prominent role in the work of Josef Matthias Hauer, whose techniques anticipated those of Schönberg. It is not unreasonable then to take the view that at the very outset of any approach to the theory of twelve-tone music the combinatoric geometry of such unordered sets should be investigated in the spirit we have pursued the Eulerian and pentatonic tonnetze.
@@
At first glance, such an assignment might seem structureless; but the permutation group for six objects is very special and the highly nontrivial structure of this group is directly related to the particular type of tone network that we wish to propose for twelve-tone music. The method that we consider is based on Sylvester’s theory of duads and synthemes (Sylvester 1844, 1861).
A duad is an unordered pair of distinct tones taken from $H_{0}$. A little thought shows there are fifteen duads. Thus, a duad $D$ is an unordered set of the form
@@
Thus, we have constructed a combinatoric configuration. More precisely, we have two sets, viz., the set of duads and the set of synthemes. Each has fifteen elements and they are in three-to-three correspondence. We say that the duad $D$ is incident with the syntheme $S$ if it is among the three duads contained in $S$. Inspection of Tables II and IV shows that two synthemes have at most one duad in common. This gives us an incidence structure and we can construct the corresponding Levi graph.
@@
Table 1: Each of the six letters corresponds to a collection of five number-synthemes called a number-total. In each number-total one finds fifteen number-duads.
@@ 66 $af,bc,de$ $ad,be,cf$ $ae,bf,cd$ $ac,bd,ef$ $ab,ce,df$
Table 3: Each of the six numbers corresponds to a collection of five non-overlapping letter-synthemes called a letter-total. In each letter-total one finds fifteen letter-duads.
@@
There is one more layer of structure, which comes as a surprise, and this is the following. Consider a collection of five synthemes with the property that each duad appears exactly once. This makes sense since there are fifteen duads and a collection of five synthemes must include fifteen duads. A collection of five ‘non-overlapping’ synthemes is called a ‘total’ (again, Sylvester’s term). There are exactly six such totals that can be constructed. Let us label the six totals with lower case Roman letters. Thus, the totals are elements of the set $\{a,b,c,d,e,f\}$. Since we have a second set of six elements, we can form the corresponding ‘duads’ and ‘synthemes’ for this set as well. We shall call these letter-duads and letter-synthemes, respectively. Thus, a letter-syntheme is an unordered triple of letter-duads. Then by a letter-total we mean a collection of five non-overlapping letter-synthemes. Two distinct totals have one syntheme in common. Hence, the letter duads of the totals are the original synthemes; and it follows that the letter-synthemes are the original duads. The relations between duads, synthemes, totals, letter-duads, letter-synthemes, and letter-totals can be seen in Tables I – IV (Richmond 1900, Baker 1925, Coxeter 1958).
@@
The advantage of using the number-duads and the letter-duads is that this allows us to label the resulting Levi graph neatly, where the number-duads, which correspond to letter-synthemes, are represented by white vertices and the letter-duads, which correspond to number-synthemes, are represented by black vertices. In summary, one sees that there is indeed a rich hierarchical structure implicit in the specification of a hexachord.
The Levi graph of the Cremona-Richmond $\{15_{3}\}$ is shown in Figure 16 (Tutte 1947, Coxeter 1950, Brier & Bryant 2021). We can take this graph as representing a tonnetz for the twelve-tone system when an unordered hexachord has been selected as a primitive. Its girth is eight: there are no hexacycles.
@@
A construction of the $\{15_{3}\}$ is as follows. Here we find it convenient to use the language of projective geometry. We begin with six points $p_{1},.\,.\,.,p_{6}$ in general position in the four-dimensional real projective space $\mathbb{RP}^{4}$. When joined pairwise, these points determine fifteen lines. Let us write $L_{12}=p_{1}\vee p_{2}$ for the line joining $p_{1}$, $p_{2}$, and $H_{3456}=\vee(p_{3}p_{4}p_{5}p_{6})$ for the complementary hyperplane joining $p_{3}$, $p_{4}$, $p_{5}$, $p_{6}$. This hyperplane is an $\mathbb{RP}^{3}$. Each of the fifteen lines meets its complementary hyperplane at a point. Thus, for example, we can write $P_{12}=L_{12}\wedge H_{3456}$ for the meet of $L_{12}$ and the hyperplane $H_{3456}$. There are evidently fifteen such ‘meeting points’. On the other hand, for example, we also have $P_{34}=L_{34}\wedge H_{1256}$ and $P_{56}=L_{56}\wedge H_{1234}$. Let us write $S_{12,34,56}$ for the line given by the meet of the three hyperplanes. Since $H_{1234}$ and $H_{1256}$ both contain $L_{12}$ it follows that the plane of intersection $D_{12}$ of these two hyperplanes contains $L_{12}$. Then since $D_{12}$ meets $H_{3456}$ at $S_{12,34,56}$ it must be that $L_{12}$ meets $S_{12,34,56}$ at $P_{12}$. By symmetry, all three of the ‘duad’ meeting points $P_{12}$, $P_{34}$, $P_{56}$ lie on the ‘syntheme’ line $S_{12,34,56}$. The fifteen meeting points are the duads of the configuration; and the fifteen lines of three-at-a-time hyperplane intersection, each of which contains three of the duads, are the synthemes of the configuration. That gives the four dimensional version of the construction, which can then be projected to a plane to provide a $\{15_{3}\}$ in two dimensions, the Cremona-Richmond configuration. Once the result has been obtained on the real projective plane, it can be Euclideanized and that takes us back to Figure 17. For further details of the construction see Richmond (1900), Baker (1925) and Coxeter (1950). Summing up the results of this section, we have the following:
@@
A tonnetz can be associated with the selection of an unordered hexachord in twelve-tone music, given by a self-dual $\{15_{3}\}$ in $\mathbb{R}^{2}$ known as the Cremona-Richmond configuration. In this tonnetz, the $15$ ‘major’ duad chords are represented by points and the $15$ ‘minor’ syntheme chords are represented by lines. The $45$ incidence relations between the $15$ points and the $15$ lines determine the edges of the corresponding Levi graph, which is known as Tutte’s $8$-cage.
The cycles of the twelve-tone tonnetz are summarized in the Appendix, Table VIII. It is interesting to note that there are 144 Hamiltonian cycles altogether. There are no hexacycles, as we said, but there are 90 octacycles. Among these are the five overlapping 1pp-octacycles, which range about the perimeter Hamiltonian in a fashion that is very similar to the arrangement of the five overlapping 1pp-hexacycles in the pentatonic tonnetz. One might not have thought a priori that there were any particular affinities between the pentatonic system and the twelve-tone system, but there are. In fact, there is a curious kind of ‘five-ness’ that pervades the twelve-tone system that is clearly evident in our approach (e.g., in Table VIII) though not perhaps so obvious in traditional treatments of serialism. In Figure 18 we present a short composition for violin based on this pentagonal symmetry.  [^22] If any one tone of the hexachord is singled out (in this case, the tone 4), the remaining tones can be arranged with pentagonal symmetry. One example of ‘five-ness’ that does manifest itself in the traditional approach to the twelve-tone system is the involutory map from the set of integers mod 12 to itself given by multiplication by 5.
Figure 16: The Levi graph of the Cremona-Richmond configuration $15_{3}$ maps to a tonnetz for the system of twelve-tone music based on an unordered hexachord. Each white vertex represents a duad (number pair) and each black vertex represents a syntheme (letter pair). Each duad belongs to three synthemes and each syntheme contains three duads.
@@
Figure 18: ‘Decacycle for Violin’, based on a 5pp-decacycle of the 12-tone tonnetz constructed from a hexachord of the form $1,2,3,4,5,6]=[F_{\sharp},G_{\sharp},C_{\sharp},D_{\sharp},E,A_{\sharp}]$. The duads and synthemes of the decacycle are given by $\langle 12,ab,56,df,13,ac,25,bf,36,cd,12\rangle$, forming the sides and vertices of a pentagon in the Cremona-Richmond configuration of Figure 17. The structure of this pentagon can also be seen in Figure 16. The relations between the letter duads and the number synthemes are shown in Table II. The note $B$ is used in this composition to ground the music.